OpenStudy (anonymous):

Find the solution to the equation 64(3 – x) = 42x. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (anonymous):

Can Someone help me?

OpenStudy (anonymous):

its supposed to be 64^(3 – x) = 4^2x.

ganeshie8 (ganeshie8):

hint- 64 = 16*4 = 4*4*4 = 4^3

ganeshie8 (ganeshie8):

does that help... ?

OpenStudy (anonymous):

I dont understand

OpenStudy (eamier):

if a^(x+z) = a^y, then x+z=y

OpenStudy (anonymous):

what?

OpenStudy (eamier):

try make 64 and 4 look alike. 64 is 4 to the power of...

OpenStudy (anonymous):

3

OpenStudy (eamier):

you mean x=3. no

OpenStudy (eamier):

\[64^{3-x}=4^{2x}\] \[4^{3(3-x)}=4^{2x}\]

OpenStudy (eamier):

so \[3(3-x)=2x\]

OpenStudy (eamier):

9-3x=2x 5x=9 \[x=\frac{9}{5}\].

OpenStudy (anonymous):

oh. Thanks for the help